Home‎ > ‎Reports‎ > ‎Crew 108‎ > ‎

day02


Crew #108 Log Book for December 04, 2011

Photos of the day






Commander's Report


Date and time: 4 December, 5pm

Written by : Charlotte Poupon, Station Leader

Yesterday night, after the MS team left the Hab, we officially declared the beginning of our simulation.

It was established than even if some cellphones find network we will not use them. But, for safety reason we will take them on EVA to be located by GPS if needed.

This morning, at 9 am, the Executive Officer, hoisted the Martian Flag up to the Hab to welcome our first day of simulation. Then we had our first daily meeting to settle the crew time table.

FOOD BACKGROUND AND STUDY

Crew tidied the kitchen and organized food for the whole rotation. Usha will be the "chef", responsible of the menu. Every day a new couple of crew member will assume the responsibility of cooking.

The safety Officer, gather the weights of the whole crew as indicated in the Food study Framework. Each crew member was assigned a code (from 108-A to 108-F) to preserve privacy in this study.

We had a dehydrated soup for lunch, all together, in a good atmosphere.

Tonight we had the initial meal : rice, chili, broccoli, cooked brownies.

EVA

Two EVAs of 3 people each was organized, 3 others assuming the HabCom.

Due to the very poor quality of transmissions, we will try to established a hand-signal language to guarantee a better communication between crew members on EVA. We'll try to make a document as reference for futures EVA & rotations.

The Executive Officer will collect ours tracks during EVA.

GENERAL MOOD

Very Good.

WEATHER

Cloudy on the morning, then sunny.
Dry.


Engineering Report

Date    4th December 2011
Time    20:00

Generators/ Electricity
Kitty
Hobbs   Where is this located?
Coolant Unknown
Oil     Unknown
Battery Voltage Unknown
Casper Battery Voltage  Unknown

Honey   not used

Fuel
Diesel Tank      1/2
Propane Tank    10%
Gasoline (gallons)      10
Oil Available (quarts)  Unknown
Notes:

Inverter/ Batteries
Outside Battery Reading Where is this located?
Heat Tapes      Where are these located?
One Meter
Amps    ????
Volts   ????
Watts   ????
kWHr    ????
Vehicles
New Blue
Used    no
Oil     not checked
Tyres   good
Odometer (miles)        not checked

ATV 1
Gasoline used (gallons)   1/10
Oil Checked     no
Tyres   good

ATV 2
Gasoline used (gallons)   1/10
Oil Checked     no
Tyres   good

ATV 3
Gasoline used (gallons)   1/10
Oil Checked     no
Tyres   good

Water
Trailer (inches from bottom)    9
Outside Potable (inches from top)       6
Inside Potable (markings)       Top Green Line

Temperature (deg F)
Min     Don't know
Current Is there an external temperature sensor?
Max     Don't know
Humidity (%)
Min     Don't know
Current 10
Max     Don't know

Observatory
radio telescope not used
Musk observatory        not used
Hab
HVAC    OK
Mars Surface Suits      OK except for loose joints on 4 and 5
Computers/Network       OK
Appliances      OK
GreenHab        Offline
Observatory     not used
Weather Data Logger
Time    18:30
Temperature (deg F)     30,2
Pressure (mb)   860
Humidity (%)    98
Description     Overcast in the morning. Sunny in the afternoon.
Upgrades        None
Notes   Will repair loose joints on suits tomorrow


Journalist Report


Date and time: 4 december 2011

Written by : Charlotte Poupon

Text:

Today was our first day on simulation. Our first SOL at the Hab. After the traditional morning meeting, we organized the food for the whole rotation.

On the afternoon, we made 2 short EVAs for everybody to practice ATV driving. The first EVA found cyanobacteria and salt deposits. The second EVA went in direction of Phobos Pick and found a nice spot with beautiful white rocks in a middle of a pink sand river. Once in the Hab, two crew members started to cook the initial meal of the food study, when suddenly Marty-the-Mouse poked a single red eye out of the cupboard…

The cooking crew members jumped on chairs, but
bravely went on cooking.
After chasing Marty-the-Mouse for a while, we trapped him downstairs under a plastic box.

The crew decided to let him live and plan to release him tomorrow during a mousitarian EVA as a sign of peace towards the Martian Mouse People.
Even though the whole crew deeply respects the mouse people, none of the crew members were willing to share a piece of the delicious brownie made for dessert.

EVA 1 Report

Date and time: December 4th, 2011

Crew Members: LeClair (Sortie Lead, Geologist), Lingappa (Biologist), Framis (Journalist)

External Operations: Airlock seal broken at 1228, Re-pressurization complete at 1410. 

Primary Objectives: 

1)      Equipment and New Technology Testing

2)      Crew – Systems Familiarizations

Secondary Objectives

1)      Test Comm Range, Strength, Quality and Reliability

2)      Initial Scientific Inquiry

 

Traverse 1:  Turtle Hills

 

Traverse 2:  Collapsing Wall

 

Traverse 3:  Bob’s Rock Garden, 200m WSW of Oyster Turn

 

Results: 

 

-Transportation Units and Environmental Suits nominal

-External Radio Communications highly unreliable

-Possible Cyanobacteria found


EVA 2 Report

Date and time: 4 December 2011

Written by: Charlotte Poupon, Crew 108, Commander

Time: 03.05 pm to 04.25 pm

Crew Members: Charlotte, Ashley and Mike

Site Location:

Transit Mode: ATV

Objectives: ATV familiarization 

Results: 
Beside the navigation we were quite surprised by the very poor quality of communication between either crew members on EVA or between EVA and
the Hab. This can be a very big safety problem.

We've decided to try to establish a non-verbal communication between crew member on EVA and basic guideline for navigation.

This will be document by the end of our rotation.

Lessons Learned: TBA as above